OpenStudy (anonymous):

The speed of train A is 8mph slower than the speed of train B. Train A travels 210 miles in the same time it takes train B to travel 250 miles. Find the speed of each train.

OpenStudy (anonymous):

r t=d, t=d/r A=42 mph, B=50 mph

OpenStudy (anonymous):

rate x time = distance time = distance/rate 250/r=210/(r-8) 250(r-8)=210r 250r - 2000 = 210r 40r = 2000 r = 50 Train A = 42 mph B = 50 mph. Sorry fr the long wait, I did it wrong twice. D: